Mass of KxFe(C2O4)y · zH2O : 5.10 g

Mass of sample : 0.195 g

Mass of FeCl3 used in preparation : 1.60 g

Molarity of standard NaOH used : 0.100

V1, volume of standard NaOH required for first equivalence point : 8.950 mL

V2, volume of standard NaOH required for second equivalence point : 17.70 mL

Answer these questions using info above:

Calculate the mass of potassium in the sample :

_______g

Calculate the percent of potassium in the sample :

________%

Calculate the volume of standard NaOH that reacted with iron :

_______ mL

Calculate the mass of iron in the sample :

_______ g

Calculate the percent of iron in the sample :

________ % Fe
Answer #1

Mass of KxFe(C2O4)y · zH2O : 5.10 g

Mass of sample : 0.195 g

Mass of FeCl3 used in preparation : 1.60 g

Molarity of standard NaOH used : 0.100

V1, volume of standard NaOH required for first equivalence point : 8.950 mL

V2, volume of standard NaOH required for second equivalence point : 17.70 mL

Answer these questions using info above:

Calculate the mass of potassium in the sample :

1st equivalence point corresponds to K+ ;

mol K+ = moles of NaOH used in titration (from V1 ) = 0.100 mol/L * 8.95 ml/1000 ml/L = 0.000895 moles

mass of K+ in sample =  0.000895 moles* 39.1 g/mol = 0.0350 g

Calculate the percent of potassium in the sample :

mass of K+ in sample / mass of sample *100

0.0350 g / 0.195 g *100 = 17.95 %

Calculate the volume of standard NaOH that reacted with iron :

V2 -V1 = 17.70 ml - 8.950 ml = 8.75 mL

Calculate the mass of iron in the sample :

three moles of hydroxide ion are required to react with each mole of iron ion in the salt.

moles of Fe = 1/3* moles of NaOH used in titration (form V2 - V1) = 0.100 mol/L * 8.75 ml/1000 ml/L /3 = 0.000292 moles

mass of Fe in sample = (moles of Fe)(55.85 grams/mole) = 0.000292 moles*55.85 g/mol = 0.0163 g

Calculate the percent of iron in the sample :

mass of Fe in sample / mass of sample *100

0.0163 g / 0.195 g *100 = 8.35 %